OpenStudy (anonymous):

If a and b are even integers and c is an odd integer, which of the following would be an odd integer? a + b + c ab - bc a x b x c b x c c (a + b)

OpenStudy (anonymous):

a + b + c

OpenStudy (anonymous):

sub in values and try for yourself? ex. a + b + c 4+4+3= 11

OpenStudy (anonymous):

but how'd u get that bro?

hero (hero):

You get it by doing exactly what they just did. Subbing in values.

OpenStudy (anonymous):

by thinking, just for instance make a and b = to 2

hero (hero):

Put there is a more general rule: Even + Odd = Odd

OpenStudy (anonymous):

and amke c 3, just solve them in your head

OpenStudy (anonymous):

make*

hero (hero):

But* not Put

OpenStudy (anonymous):

Ok let a =2x b =2y c= 2m+1 Now tthe sum a+b+c must be written in the form 2n+1 Thus a+b+c =2x+2y+2m+1 =2(x+y+m)+1 Now let x+y+m =n =2n+1 W/c is odd.